# HF radar estimation of ocean wave parameters: second-order Doppler spectrum versus Bragg wave modulation approach

### v2

@@ Abstract
We propose an original technique for the HF radar estimation of the main sea state parameters by exploiting the amplitude modulation of the radar signal time series. While the classical method for ocean wave measurement is based on the second-order ocean Doppler spectrum, this alternative approach uses the slow amplitude modulation of the Bragg wave in the radar signal, which is more robust to noise than the second-order echo. We apply this method to an annual set of HF radar data in the vicinity of Toulon (Mediterranean coast of France) and compare it with the classical second-order method, using a nearby buoy and the WWIII model as ground truth. The Doppler spectrum-based method is found to be more accurate in calculating the significant wave height and the peak wave frequency while the modulation approach provides coarser estimates but can achieve longer ranges and higher temporal coverage.
@@ 1 Introduction
Figure 1: Location of HF radar coverage a) in the western Mediterranean Sea and b) zooming in on the area. The red dot marks the location of Fort Peyras, the blue dot that of the CANDHIS buoy, and the white lines represent 5 range units (every $7.5km$) and 5 azimuth units (every $5\rm{{}^{o}}$) of the radar coverage. The background color is the bathymetry measured in $m$. Toponyms cited in the main text appear in the figure.
@@
High-frequency radars (HFR) have been used for about half a century for the monitoring of coastal surface currents (Barrick et al., (1977)). As a secondary geophysical product, HFR also provide measurement of sea state parameters and wave spectra (e.g. Wyatt, (2021) for a recent review). However, this type of estimation is less accurate and less straightforward than the surface current estimation, due to theoretical as well as practical reasons. On the theoretical level, while the radial surface current only requires the measurement of the dominant first-order Bragg peaks, the wave spectrum is related to the side-band continuous components of the ocean Doppler spectrum surrounding the Bragg peaks. This second-order contribution has a complex analytical expression in the form of a nonlinear integral equation involving a coupling coefficient and quadratic combinations of the wave spectra at pairs of frequencies satisfying a resonance condition Hasselmann, (1971); Barrick, (1972); Lipa and Barrick, (1986). Under some restrictive assumptions, Barrick, 1977b ; Barrick, 1977a derived in $1977$ a method to extract the main sea state parameters from this second-order integral and accordingly, from the measured ocean Doppler spectrum. He showed that the significant wave height (SWH) and the centroïd wave frequency can be expressed as simple ratios of weighted integrals involving the first- and second-order Doppler spectra. Soon after, linear and nonlinear inversion methods for retrieving the ocean wave spectrum from single or dual radar sites were developed in a series of papers (e.g. Lipa, (1977); Wyatt, (1990); Hisaki, (1996)). Today, the inversion of the waves has reached a mature state and there are well-established techniques and commercial software for the determination of the ocean wave directional spectrum (see e.g. Wyatt, (2021) for a recent review). However, these inversion techniques require more expertise than that required to calculate the main wave parameters using Barrick’s formulas. They have also physical as well as practical limitations (Wyatt, (2000)) that restrict their range of applicability. The most critical issue relates to the quality of the radar data and, in particular, to the ability to measure the second-order Doppler spectrum with a sufficient signal-to-noise ratio. Except for the high sea state and the highest radar frequencies, the second-order Doppler spectrum is usually several tens of dB below the level of the first-order peaks and reaches sooner the noise threshold when the range is increased. Hence, wave estimation methods based on the second-order Doppler analysis are confined to a restricted radar coverage in the near to middle range, depending on the radar frequency. Another issue is the clear identification of the second-order contribution and its separation from the first-order Bragg components, which can be ambiguous especially in the case of strong currents or in the absence of azimuthal resolution. For these reasons, many authors have only considered the estimation of the bulk wave parameters using empirical methods (e.g. Essen et al., (1999); Gurgel et al., (2006); Lopez et al., (2016)) or empirical adjustments of Barrick’s method Heron and Heron, 1998a ; Maresca Jr and Georges, (1980); Ramos et al., (2009); Alattabi et al., (2019); Al-Attabi et al., (2021); Basañez et al., (2020). Most of these methods necessitate the calculation of Doppler spectra that are resolved both in range and azimuth and can thus be associated with a reduced sea surface patch with constant sea state parameters. In the case of compact antenna arrays, the Doppler spectra are nondirectional, that is, they cannot be resolved in azimuth. This does not allow for the measurement of the wave frequency spectrum from a single station. Nevertheless, the $H_{s}$ can be estimated on range arcs adapting Barrick’s original methods. The long-term performances of such measurements have been assessed by comparison with buoys and models and are quite variable depending on sea state, range, and direction of waves (see e.g. Roarty et al., (2015); Orasi et al., (2018); Saviano et al., (2020) or Lorente et al., (2022) for a recent review).
@@ 2.1 High Frequency Radar data
In this study, we use the monostatic signals collected by the Fort Peyras site (Figure 1) during $2020$, a year for which data availability was over $91\%$. This radar operates in a $16.15$ MHz center frequency band with an assigned bandwidth of $100$ kHz. Depending on weather conditions, it can cover a range of $60$ km with a range resolution of $1.5$ km and an angular sector of about $110\rm{{}^{o}}$. The data consist of the complex I/Q voltage recorded on the set of $12$ receive antennas and sorted in range after a preliminary Doppler processing in the fast time. The sampling rate in the slow time is $0.26$ seconds and the time series are acquired in continuous blocks of $4096$ values (lasting about $18$ min) every $20$ min, which corresponds to the standard “SORT file” output format of WERA HF radar. The data are further processed in azimuth by applying the classical beamforming algorithm (with steering steps of one degree) to the $12$ channels of the antenna array. This leads to a complex discrete time series $s(t)$ for each radar cell of size $1.5$ km×$\times 1$ degree and each $20$ min sequence. In the following, each radar cell is labeled by a range number ($1$ to $64$ starting from the radar) and an azimuth number ($1$ to $110$ starting from the westernmost corner).
@@ 2.2 Wave data
During the same year, the results of the Wavewatch III model (WWIII) at an hourly time resolution were available over the study area on a 3-level nested grid (with a spatial resolution of $800$m in the radar aera) forced by $10$-m wind fields obtained from the non-hydrostatic mesoscale Weather Research and Forecast model (for more details consult Lira-Loarca et al.,, 2022). The SWH and the mean wave frequency are estimated through the spectral moments (1) and the peak wave frequency $f_{p}$ is also provided. The annual mean values of the SWH and the peak wave frequency in the coastal region of Toulon according to the WWIII model are shown in Figure 2. As seen, the typical SWH in the radar aera is of the order of $1$ m while the typical value of the peak frequency is about $0.22$ s. While the ground truth in the vicinity of the radar area is limited to the single location of the CANDHIS buoy, it is possible to use the WWIII model to complement the analysis. To check the accuracy of the WWIII in the radar area, we compared the time series of $H_{s}$ and $f_{p}$ at the buoy location to the actual measurements. The point by point comparison is shown in the scatterplots of Figures 3-4. The Root Mean Square Difference (RMSD) of SWH is $25$ cm, which is about $25\%$ of its mean value ($0.96$ m) while the RMSD of the peak wave frequency is $0.037$, which is about $20\%$ of its mean value ($0.19$ Hz). The Pearson correlation coefficient is higher for the SWH ($0.94$) than for the peak wave frequency ($0.70$). The statistical parameters to compare the different time series are summarized in Table 1.
@@ 3 Material and Methods
Two different methods have been used to estimate the ocean waves, $H_{s}$ and $f_{p}$, from HF radar signals. The first is the classical Second Order Doppler Spectrum method, first introduced by Barrick, 1977a ; Barrick, 1977b , while the second is an original method that uses only the time modulation of Bragg wave amplitude. In the following we will refer to these methods simply as “Barrick’s method” and the “Modulation method”.
@@ 3.1 Second Order Doppler Spectra Method
The classical method for estimating $H_{s}$ and $f_{p}$ from the HF sea echo originates from the seminal work of Barrick, 1977a ; Barrick, 1977b and involves straightforward calculations based on the ocean Doppler spectrum. The hourly range-azimuth resolved Doppler spectra are classically obtained by forming the power spectrum of the radar signal:
@@
where $\hat{s}(f)$ is computed in the standard way on a discrete set of frequencies using a Fast Fourier Transform of the time series with a Hamming attenuation window and an integration time of about $2$ min. To reduce the noise variance, a further incoherent summation is applied by taking a median average of the sample spectra $\lvert\hat{s}(f)\rvert^{2}$ over $15$ overlapping intervals within the $18$ min duration of each available SORT file (3 per hour). This will result in an average Doppler spectrum with a bandwidth of $4$ Hz (corresponding to a sampling rate of $0.26$ sec) and a frequency resolution of $\Delta f=1/T\simeq 0.008$ Hz. An example of such a Doppler spectrum is shown in Figure 5, a) for the $8$ range cell ($12$ km) and azimuth $40$ (marked with orange lines in Figure 1, b) for March $8$, $2020$. The corresponding range Doppler spectrum along the same direction is shown in Figure 6. In this example, the strongest first-order Bragg line extends to about $90$ km, while the strongest second-order sideband is visible up to $65$ km away from the radar. The dashed white circles highlight a power grid artifact ($50$ Hz harmonics) that contaminates the low Doppler frequency content of the radar signal at some ranges; this electrical artifact is particularly pronounced at the Fort Peyras site due to the imperfect power supply at this location. It does not affect the estimation of the radial current from the first-order Bragg lines but must be treated with caution when dealing with the full ocean Doppler spectrum.
@@
The main challenge of Barrick’s method is to effectively discriminate between first and second-order spectra and to accurately distinguish signal from noise. To overcome this, specific techniques have been developed that allow the precise setting of separation thresholds and the establishment of control criteria for integration regions. The effectiveness of this technique is limited by its dependence on the second-order Doppler spectrum, which restricts its application to shorter distances where this spectrum remains well above the noise level. Due to this greater sensitivity to noise, it is important to subtract the mean noise level from the second-order Doppler spectrum before performing the integration in (5) and (7).
To determine the exact point of transition from first to second-order spectra, and thus avoid false boundaries, the following method is used: positive peaks (marked with a blue circle in Figure 5, b) and negative peaks (marked with red circles in Figure 5, b) are selected. The minimum negative peak (marked with a solid red circle in Figure 5, b) and the maximum positive peak at frequencies higher than the minimum negative peak (marked with a solid blue circle in Figure 5, b) are identified. We impose the condition that the intensity difference between the first-order peak and the minimum negative peak ($d_{1}$) must be at least twice the difference between the minimum negative peak and the maximum positive peak ($d_{2}$). If this condition is not met, the selected minimum negative peak is discarded and the process is repeated with another negative peak initially identified. This threshold value of $2$ is purely empirical and has been found to be effective in removing satellite peaks near the first-order Bragg peak that could be misinterpreted as second-order peaks. Although these peaks are located at very close frequencies, the resulting error in calculating the integral of the second-order Doppler spectrum can be large, resulting in an inaccurate estimate of $H_{s}$.
@@ 3.2 The Bragg Wave Modulation Method
where $t$ is time; $A^{\pm}$ are complex amplitudes of the form (8) with random phases; $\tilde{f}_{B}$ is the modulated Bragg frequency and $f_{u}=2u/\lambda$ is the additional frequency shift due to the radial surface current $u$. By hypothesis, the variance of the amplitude is proportional to the variance of the long wave steepness:
$\mathrm{\langle\lvert A_{B}^{+}\rvert^{2}\rangle}\sim\mathrm{\langle\lvert k_{p}a_{p}^{\pm}\rvert^{2}\rangle}$ (11)
@@
$H_{s}^{2}\sim\mathrm{\langle\lvert A_{B}^{\pm}\rvert^{2}\rangle}$ (12)
The complex Bragg wave amplitudes $A^{\pm}$ are generally different but if we assume that they are uncorrelated due to random phase shifts between the approaching and receding waves, we can simply sum their contribution to the SWH:
$H_{s}^{2}\sim\mathrm{\langle\lvert{A_{B}^{+}}\rvert^{2}\rangle}+\mathrm{\langle\lvert{A_{B}^{-}}\rvert^{2}\rangle}$ (13)
@@ 4.1 Statistical performances at the buoy location
The CANDHIS buoy is located at a distance of about 36 km from the radar. Although this relatively large distance is not optimal for estimating the wave parameters from the HFR data, the buoy data can be used to assess the performance of the two proposed methods. Figures 8 and 9 show the time series (re-centered over one month for better visibility) for the SWH and peak wave frequency, respectively, using either the HFR-based estimates, the CANDHIS records or the WWIII model. It appears that both methods give very consistent estimates compared to the buoy measurements, with Barrick’s method appearing to be more accurate but also more lacunary than the Bragg wave modulation method. The plots also suggest that the WWIII estimate is often less accurate than the radar measurements, a fact which is confirmed by a thorough statistical analysis (see further).
To compare the methods in more detail, the radar/buoy scatterplots are shown in Figures 10 and 11 for the annual time series where simultaneous data are available. Along with these plots we calculated the usual statistical parameters, namely the RMSD and the Pearson correlation coefficient (R) (see Table 2). It confirms a better correlation of Barrick’s method with the buoy measurements ($R=0.87$ and $R=0.62$ for the SWH and peak wave frequency, respectively ) while the estimates obtained with the Bragg Wave Modulation Method are more scattered ($R=0.43$ and $R=0.38$, respectively). It also shows that Barrick’s method has a tendency to overestimate both $H_{s}$ and $f_{p}$. This justifies the empirical scaling factor which has been used by several authors (Maresca Jr and Georges,, 1980; Heron and Heron, 1998b, ; Ramos et al.,, 2009) to adjust the radar estimates of $H_{s}$. We found that a scaling factor of about $0.58$ significantly improves the agreement with the buoy and reduces the RMSD from $0.48$ m to $0.29$ m, which is close to the RMSD obtained with the WWIII model ($0.25$ m). Note that this scaling factor is very consistent with the values found in (Maresca Jr and Georges,, 1980) ($0.55$) Ramos et al., (2009) ($0.58$).
Figure 8: Time series of $H_{s}$ (m) from February 1st to April 30, 2020 at the location of the CANDHIS buoy, as estimated by the WWIII model (red lines), Barrick’s method (blue symbols) or the Modulation method (magenta symbols). The buoy measurements are shown with black solid lines.
@@
Table 2: Mean bias ($\mu$), RMSD ($\sigma$) and Pierson correlation coefficient ($R$) of Barrick’s method (B), the Modulation method (M) and the WWIII model (W) for the estimation of $H_{s}$ (in meter) and peak wave frequency $f_{p}$ (in Hz) over the year $2020$ by comparison with the CANDHIS buoy measurements (range number $21$ and azimuth number $78$). The RMSD of $H_{s}$ following from Barrick’s estimate is given with ($0.48$ m) and without ($0.29$ m) the scaling factor $0.58$.

### v3

@@ Abstract
We propose an original technique for the HF radar estimation of the main sea state parameters by exploiting the amplitude modulation of the radar signal time series. While the classical method for ocean wave measurement is based on the second-order ocean Doppler spectrum, this alternative approach uses the slow amplitude modulation of the Bragg wave in the radar signal. Using a nearby buoy and the WWIII model as ground truth, we apply this method to an annual set of HF radar data in the vicinity of Toulon (Mediterranean coast of France) and compare it with the classical formulas proposed by Barrick (1977) to derive the main wave parameters from the Doppler spectrum. Although the modulation approach is physically attractive and easier to implement, we find that the Doppler spectrum-based method is far superior at this stage in calculating the significant wave height and the peak wave frequency, even at long range.
@@ 1 Introduction
High-frequency radars (HFR) have been used for about half a century for the monitoring of coastal surface currents (Barrick et al., (1977)). As a secondary geophysical product, HFR also provide measurement of sea state parameters and wave spectra (e.g. Wyatt, (2021) for a recent review). However, this type of estimation is less accurate and less straightforward than the surface current estimation, due to theoretical as well as practical reasons. On the theoretical level, while the radial surface current only requires the measurement of the dominant first-order Bragg peaks, the wave spectrum is related to the side-band continuous components of the ocean Doppler spectrum surrounding the Bragg peaks. This second-order contribution has a complex analytical expression in the form of a nonlinear integral equation involving a coupling coefficient and quadratic combinations of the wave spectra at pairs of frequencies satisfying a resonance condition Hasselmann, (1971); Barrick, (1972); Lipa and Barrick, (1986). Under some restrictive assumptions, Barrick, 1977b ; Barrick, 1977a derived in $1977$ a method to extract the main sea state parameters from this second-order integral and accordingly, from the measured ocean Doppler spectrum. He showed that the significant wave height (SWH) and the centroïd wave frequency can be expressed as simple ratios of weighted integrals involving the first- and second-order Doppler spectra. Soon after, linear and nonlinear inversion methods for retrieving the ocean wave spectrum from single or dual radar sites were developed in a series of papers (e.g. Lipa, (1977); Wyatt, (1990); Hisaki, (1996)). Today, the inversion of the waves has reached a mature state and there are well-established techniques and commercial software for the determination of the ocean wave directional spectrum (see e.g. Wyatt, (2021) for a recent review). However, these inversion techniques require more expertise than that required to calculate the main wave parameters using Barrick’s formulas. They have also physical as well as practical limitations (Wyatt, (2000)) that restrict their range of applicability. The most critical issue relates to the quality of the radar data and, in particular, to the ability to measure the second-order Doppler spectrum with a sufficient signal-to-noise ratio. Except for the high sea state and the highest radar frequencies, the second-order Doppler spectrum is usually several tens of dB below the level of the first-order peaks and reaches sooner the noise threshold when the range is increased. Hence, wave estimation methods based on the second-order Doppler analysis are confined to a restricted radar coverage in the near to middle range, depending on the radar frequency. Another issue is the clear identification of the second-order contribution and its separation from the first-order Bragg components, which can be ambiguous especially in the case of strong currents or in the absence of azimuthal resolution. For these reasons, many authors have only considered the estimation of the bulk wave parameters using empirical methods (e.g. Essen et al., (1999); Gurgel et al., (2006); Lopez et al., (2016)) or empirical adjustments of Barrick’s method Heron and Heron, 1998a ; Maresca Jr and Georges, (1980); Ramos et al., (2009); Alattabi et al., (2019); Al-Attabi et al., (2021). Most of these methods necessitate the calculation of Doppler spectra that are resolved both in range and azimuth and can thus be associated with a reduced sea surface patch with constant sea state parameters. In the case of compact antenna arrays, the Doppler spectra are nondirectional, that is, they cannot be resolved in azimuth. This does not allow for the measurement of the wave frequency spectrum from a single station. Nevertheless, the $H_{s}$ can be estimated on range arcs adapting Barrick’s original methods. The long-term performances of such measurements have been assessed by comparison with buoys and models and are quite variable depending on sea state, range, and direction of waves (see e.g. Roarty et al., (2015); Orasi et al., (2018); Saviano et al., (2020) or Lorente et al., (2022) for a recent review).
@@
Figure 1: Location of HFR coverage in the Mediterranean region of Toulon (France). The red dot marks the location of Fort Peyras, the blue dot that of the CANDHIS buoy, and the white lines represent 5 range units (every 7.5 km) and 5 azimuth units (every $5\rm{{}^{o}}$) of the radar coverage.
@@ 2.1 High Frequency Radar data
In this study, we use the monostatic signals collected by the Fort Peyras site (figure 1) during $2020$, a year for which data availability was over $91\%$. This radar operates in a $16.15$ MHz center frequency band with an assigned bandwidth of $100$ kHz. Depending on weather conditions, it can cover a range of $60$ km with a range resolution of $1.5$ km and an angular sector of about $110\rm{{}^{o}}$. The data consist of the complex I/Q voltage recorded on the set of $12$ receive antennas and sorted in range after a preliminary Doppler processing in the fast time. The sampling rate in the slow time is $0.26$ seconds and the time series are acquired in continuous blocks of $4096$ values (lasting about $18$ min) every $20$ min, which corresponds to the standard “SORT file” output format of WERA HFRr. The data are further processed in azimuth by applying the classical beamforming algorithm (with steering steps of one degree) to the $12$ channels of the antenna array. This leads to a complex discrete time series $s(t)$ for each radar cell of size $1.5$ km×$\times 1$ degree and each $20$ min sequence. In the following, each radar cell is labeled by a range number ($1$ to $64$ starting from the radar) and an azimuth number ($1$ to $110$ starting from the westernmost corner).
@@ 2.2 Wave data
During the same year, the results of the Wavewatch III model (WWIII) at an hourly time resolution were available over the study area on a 3-level nested grid (with a spatial resolution of $800$m in the radar area) forced by $10$-m wind fields obtained from the non-hydrostatic mesoscale Weather Research and Forecast model (for more details consult Lira-Loarca et al.,, 2022). The SWH and the mean wave frequency are estimated through the spectral moments (1) and the peak wave frequency $f_{p}$ is also provided. The annual mean values of the SWH and the peak wave frequency in the coastal region of Toulon according to the WWIII model are shown in figure 2. As seen, the typical SWH in the radar aera is of the order of $1$ m while the typical value of the peak frequency is about $0.22$ s. While the ground truth in the vicinity of the radar area is limited to the single location of the CANDHIS buoy, it is possible to use the WWIII model to complement the analysis. To check the accuracy of the WWIII in the radar area, we compared the time series of $H_{s}$ and $f_{p}$ at the buoy location to the actual measurements. The point by point comparison is shown in the scatterplots of figures 3-4. The Root Mean Square Difference (RMSD) of SWH is $25$ cm, which is about $25\%$ of its mean value ($0.96$ m) while the RMSD of the peak wave frequency is $0.037$, which is about $20\%$ of its mean value ($0.19$ Hz). The Pearson correlation coefficient is higher for the SWH ($0.94$) than for the peak wave frequency ($0.70$). The statistical parameters to compare the different time series are summarized in Table 1.
@@ 3 Material and Methods
Two different methods have been used to estimate the ocean waves, $H_{s}$ and $f_{p}$, from HFR signals. The first is the classical Second Order Doppler Spectrum method, first introduced by Barrick, 1977a ; Barrick, 1977b , while the second is an original method that uses only the time modulation of Bragg wave amplitude. In the following we will refer to these methods simply as “Barrick’s method” and the “Modulation method”.
@@ 3.1 Second Order Doppler Spectra Method
The classical method for estimating $H_{s}$ and $f_{p}$ from the HFR sea echo originates from the seminal work of Barrick, 1977a ; Barrick, 1977b and involves straightforward calculations based on the ocean Doppler spectrum. The hourly range-azimuth resolved Doppler spectra are classically obtained by forming the power spectrum of the radar signal:
@@
where $\hat{s}(f)$ is computed in the standard way on a discrete set of frequencies using a Fast Fourier Transform of the time series with a Hamming attenuation window and an integration time of about $2$ min. To reduce the noise variance, a further incoherent summation is applied by taking a median average of the sample spectra $\lvert\hat{s}(f)\rvert^{2}$ over $15$ overlapping intervals within the $18$ min duration of each available SORT file (3 per hour). This will result in an average Doppler spectrum with a bandwidth of $4$ Hz (corresponding to a sampling rate of $0.26$ sec) and a frequency resolution of $\Delta f=1/T\simeq 0.008$ Hz. An example of such a Doppler spectrum is shown in figure 5, a) for the $8$ range cell ($12$ km) and azimuth $40$ for March $8$, $2020$. The corresponding range Doppler spectrum along the same direction is shown in figure 6. In this example, the strongest first-order Bragg line extends to about $90$ km, while the strongest second-order sideband is visible up to $65$ km away from the radar. The dashed white circles highlight a power grid artifact ($50$ Hz harmonics) that contaminates the low Doppler frequency content of the radar signal at some ranges; this electrical artifact is particularly pronounced at the Fort Peyras site due to the imperfect power supply at this location. It does not affect the estimation of the radial current from the first-order Bragg lines but must be treated with caution when dealing with the full ocean Doppler spectrum.
@@
The main challenge of Barrick’s method is to effectively discriminate between first and second-order spectra and to accurately distinguish signal from noise. To overcome this, specific techniques have been developed that allow the precise setting of separation thresholds and the establishment of control criteria for integration regions. The effectiveness of this technique is limited by its dependence on the second-order Doppler spectrum, which restricts its application to shorter distances where this spectrum remains well above the noise level. Due to this greater sensitivity to noise, it is important to subtract the mean noise level from the second-order Doppler spectrum before performing the integration in (5) and (7). The first-order Bragg peaks were identified in a classical way by seaching the maxima of the Doppler spectrum inside the first-order Bragg region, whose maximal excursion is determined by the maximal value of the radial current in the area. To determine the exact point of transition from first to second-order spectra, and thus avoid false boundaries, the following method is used: positive peaks (marked with a blue circle in figure 5, b) and negative peaks (marked with red circles in figure 5, b) are selected. The minimum negative peak (marked with a solid red circle in figure 5, b) and the maximum positive peak at frequencies higher than the minimum negative peak (marked with a solid blue circle in figure 5, b) are identified. We impose the condition that the intensity difference between the first-order peak and the minimum negative peak ($d_{1}$) must be at least twice the difference between the minimum negative peak and the maximum positive peak ($d_{2}$). If this condition is not met, the selected minimum negative peak is discarded and the process is repeated with another negative peak initially identified. This threshold value of $2$ is purely empirical and has been found to be effective in removing satellite peaks near the first-order Bragg peak that could be misinterpreted as second-order peaks. Although these peaks are located at very close frequencies, the resulting error in calculating the integral of the second-order Doppler spectrum can be large, resulting in an inaccurate estimate of $H_{s}$.
@@ 3.2 The Bragg Wave Modulation Method
where $t$ is time; $A^{\pm}_{B}$ are complex amplitudes of the form (8) with random phases; $\tilde{f}_{B}$ is the modulated Bragg frequency and $f_{u}=2u/\lambda$ is the additional frequency shift due to the radial surface current $u$. By hypothesis, the variance of the amplitude is proportional to the variance of the long wave steepness:
$\mathrm{Var}{(\lvert A_{B}^{\pm}\rvert)}\sim\mathrm{Var}{(k_{p}a_{p}^{\pm})}$ (11)
@@
$H_{s}^{2}\sim\mathrm{Var}{(\lvert A_{B}^{\pm}\rvert)}$ (12)
The complex Bragg wave amplitudes $A^{\pm}_{B}$ are generally different but if we assume that they are uncorrelated due to random phase shifts between the approaching and receding waves, we can simply sum their contribution to the SWH:
$H_{s}^{2}\sim\mathrm{Var}{(\lvert A_{B}^{+}\rvert)}+\mathrm{Var}{(\lvert A_{B}^{-}\rvert)}$ (13)
@@ 4.1 Statistical performances at the buoy location
The CANDHIS buoy is located at a distance of about 32 km from the radar. Although this relatively large distance is not optimal for estimating the wave parameters from the HFR data, the buoy data can be used to assess the performance of the two proposed methods. Figures 8 and 9 show the time series (re-centered over one month for better visibility) for the SWH and peak wave frequency, respectively, using either the HFR-based estimates, the CANDHIS records or the WWIII model. It appears that Barrick’s method gives very consistent estimates compared to the buoy measurements after correction of the $H_{s}$ estimate with an overall $0.58$ scaling factor. The Bragg wave modulation method, however, is less reliable in estimating the wave parameters, althrough its estimates clearly follow the observed trends of variations.
To compare the methods in more detail, the radar/buoy scatterplots are shown in figures 10 and 11 for the annual time series where simultaneous data are available. Along with these plots we calculated the usual statistical parameters, namely the RMSD and the Pearson correlation coefficient (R) (see Table 2). It confirms a better correlation of Barrick’s method with the buoy measurements ($R=0.87$ and $R=0.62$ for the SWH and peak wave frequency, respectively ) while the estimates obtained with the Bragg Wave Modulation Method are more scattered ($R=0.43$ and $R=0.38$, respectively). It also shows that Barrick’s method has a tendency to overestimate both $H_{s}$ and $f_{p}$, calling for an extra corrective factor. Such empirical scaling factor has been used by several authors (Maresca Jr and Georges,, 1980; Heron and Heron, 1998b, ; Ramos et al.,, 2009) to adjust the radar estimates of $H_{s}$. One invoked reason for applying such correction is that the weighting function used in Barrick’s formulas (5) and (7) favors waves that propagate perpendicular to the beam direction Wyatt, (2021). We found that a scaling factor of about $0.58$ significantly improves the agreement with the buoy and reduces the RMSD from $0.48$ m to $0.29$ m, which is close to the RMSD obtained with the WWIII model ($0.25$ m). Note that this scaling factor is very consistent with the values found in (Maresca Jr and Georges,, 1980) ($0.55$) Ramos et al., (2009) ($0.58$).
Figure 8: Time series of $H_{s}$ (m) from February 1st to April 30, 2020 at the location of the CANDHIS buoy, as estimated by Barrick’s method (blue lines) or the Modulation method (magenta lines). The buoy measurements are shown with black solid lines.
@@
Table 2: Mean bias ($\mu$), RMSD ($\sigma$) and Pearson correlation coefficient ($R$) of Barrick’s method (B), the Modulation method (M) and the WWIII model (W) for the estimation of $H_{s}$ (in meter) and peak wave frequency $f_{p}$ (in Hz) over the year $2020$ by comparison with the CANDHIS buoy measurements (range number $21$ and azimuth number $78$). The RMSD of $H_{s}$ following from Barrick’s estimate is given without ($0.48$ m) and with ($0.29$ m) the scaling factor $0.58$.